# UDKAG: Augmenting Large Vision-Language Models with Up-to-Date Knowledge

### v1

@@ Abstract
Large vision-language models (LVLMs) are ignorant of the up-to-date knowledge, such as LLaVA series, because they cannot be updated frequently due to the large amount of resources required, and therefore fail in many cases. For example, if a LVLM was released on January 2024, and it wouldn’t know the detailed plot of the new movie Dune 2, which wasn’t released until February 2024. To solve the problem, a promising solution is to provide LVLMs with up-to-date knowledge via internet search during inference, i.e., internet-augmented generation (IAG), which is already integrated in some closed-source commercial LVLMs such as GPT-4V. However, the specific mechanics underpinning them remain a mystery. In this paper, we propose a plug-and-play framework, for augmenting existing LVLMs in handling visual question answering (VQA) about up-to-date knowledge, dubbed UDKAG. A hierarchical filtering model is trained to effectively and efficiently find the most helpful content from the websites returned by a search engine to prompt LVLMs with up-to-date knowledge. To train the model and evaluate our framework’s performance, we propose a pipeline to automatically generate news-related VQA samples to construct a dataset, dubbed UDK-VQA. A multi-model voting mechanism is introduced to label the usefulness of website/content for VQA samples to construct the training set. Experimental results demonstrate the effectiveness of our framework, outperforming GPT-4V by $\sim$25in accuracy.
@@ 1 Introduction
Large vision-language models (LVLMs, e.g., GPT-4V [19], Gemini Series [20], and Grok [21]) have received much attention for their impressive generative capabilities. They require a large resource for data collection, cleaning, and training, restricting them from frequently updating models. However, new information and knowledge are created every time, making LVLMs ineffective in many scenarios. For example, if we talk with LLaVA-1.6 [23] (released on January 30, 2024) about the detailed plot of the new movie Dune 2, such as “the dead of an important duel", it performs very badly. It is promising to augment LVLMs by retrieving up-to-date knowledge via internet search during inference, i.e., internet-augmented generation (IAG). Although commercial LVLMs such as GPT-4V [19] and Claude3 [22] have the ability of IAG, the specific mechanics underpinning them remain undisclosed. This paper proposes a plug-and-play framework to augment different LVLMs in handling visual question answering (VQA) about up-to-date knowledge, named UDKAG.
@@
Figure 1: The proposed UDKAG, a framework for LVLMs to access up-to-date knowledge.
@@
To validate the effectiveness and generalizability of the proposed framework, we incorporate 13 state-of-the-art LVLMs into the framework, such as GPT-4V [19] and LLaVA-1.6 [23]. Notably, once the hierarchical filtering model is trained, our framework can adapt different LVLMs and improve their performance without any fine-tuning. Extensive experimental results demonstrate that our framework can significantly improve LVLM’s ability to answer questions about up-to-date knowledge. Incorporating the LLaVA-1.6 model of our framework even outperforms the self-contained IAG-capable GPT-4V by $\sim$25in accuracy on UDK-VQA test set.
The contributions of our work are summarized as follows. (1) We propose the first open-source framework which seamlessly incorporates existing LVLMs to equip them with up-to-date knowledge during inference. (2) We propose a pipeline which can automatically generate VQA samples related to up-to-date news, and construct the first test set for evaluating the ability of LVLMs in handling VQA about up-to-date knowledge. (3) Extensive experimental results on 13 state-of-the-art LVLMs demonstrate the effectiveness of our framework.
@@ 2.2 Large Models with Search Engine
3 Framework
In this section, we introduce UDKAG, a framework that seamlessly incorporate existing LVLMs, allowing these LVLMs to access up-to-date knowledge without fine-tuning. The whole framework is illustrated in Figure 1. For a natural language question $Q$ about an image $V$, we first extract queries for both $Q$ and $V$ via the query generator. Then we enter the queries into search engines, and the search engine would return related websites, each of which consist of a title and a snippet. To identify the most helpful content within the websites, a website filter is used to filter the websites based on their titles and snippets, and a content filter is further used to filter the content of the websites filtered by the website filter. Finally, we stitch the filtered content together to prompt existing LVLMs.
@@ 3.1 Query Generator
Question Query Generator. To get queries that make search engines return websites containing helpful content, we leverage large language models (LLMs) to extract queries for $Q$. Thanks to the language understanding capability of LLMs, the role played by each word can be well inferred from the grammatical information of $Q$, even if certain words are unknown for the LLMs. We use “Do not try to answer the question, just print the most informative no more than three entities in the question. Put them on one line and separate them with comm." to prompt LLMs to generate queries.
@@ 3.2 Search Engine
The extracted queries are fed into a search engine, and the search engine returns relevant websites with their titles and snippets. However, the returned titles and snippets often contain limited and incomplete information. For example, for a website with title “Pororo Dragon Castle Adventure", the entire snippet returned by Bing is “Pororo and his friends were having fun when a little red dragon named Arthur appears above them Arthur who claims to be the king of dragons commands Pororo and his friends to search for his Dragon …", obviously there is more about “Pororo Dragon Castle Adventure" contained in the website. Thus we parse the textual content of all websites. For a website, not all of its content contributes to answering questions, we empirically divide the website content into segments every third sentence for a more granular selection of content.
@@ 3.3 Hierarchical Filtering Model
Website Filter. The aim of the website filter is to perform the filtering of websites based on their titles and snippets. Specially, a website scoring model is trained via instruction tuning, to predict how helpful a website will be in answering a question, and the $N$ websites with higher scores would be kept. The training samples are in the format $(T,S,Q,V,R_{w})$, where $R_{w}$ is a quantitative usefulness in the interval $[0,1]$ representing how helpful a website with title $T$ and snippet $S$ will be in answering $Q$ related to $V$. Based on the samples, we construct instructions like “How helpful is an article with such a title and snippet in answering the question based on the image? Choose the best option. Title: <$T$> Snippet: <$S$> Question: <$Q$> Options: A. 1.0 B. 0.8 C. 0.6 D. 0.4 E. 0.2 F. 0.0". In doing so, the score regression problem is converted into a classification problem, which is easier to learn.
Content Filter. The content filter is used to select the most helpful content segments from the websites filtered by the website filter. For each content segment, we predict how helpful is it for answering $Q$ by a content scoring model. The content scoring model is trained by samples in the format $(C,Q,V,R_{c})$, where $C$ is a content segment, and $R_{c}$ is the quantitative usefulness of $C$ in answering $Q$. The instructions for training the content scoring model are in the format: “How helpful is this context in answering the question based on the image? Choose the best option. Context: <$C$> Question: <$Q$> Options: A. 1.0 B. 0.8 C. 0.6 D. 0.4 E. 0.2 F. 0.0". We use the model to sort all content segments and select the $M$ highest scoring ones as the obtained segments.
@@ 3.4 Augmented Generation
We augment existing LVLMs by prompting them with the final obtained content, to improve their ability of answering questions about up-to-date knowledge. Taking answer the multiple choice questions as an example, for a question $Q$ with candidate answers $A_{1}$, $A_{2}$, $A_{3}$ and $A_{4}$, we use the prompt “Given context: <$X$> Question: <$Q$> Answers: A.<$A_{1}$> B.<$A_{2}$> C.<$A_{3}$> D.<$A_{4}$> Answer with the option’s letter from the given choices directly based on the context and the image.", where $X$ denotes the final content obtained by the hierarchical filtering model.
@@ 4.2 Question Generation
For each query, we use Bing to search for relevant and up-to-date news. For the scraped news content, we divide every third sentence into a segment, and use the following message to prompt GPT-3.5 to generate a question-answer pair and several confused answers for each segment: “Given context: <$Con$> Filling the blanks to generate a question about the most informative event of the context, generate an correct answer to the question in no more than three words based on context, and generate three incorrectly confused answers of no more than three words based on context. Question:       Correct answer:       Incorrect answers: A.       B.       C.      ", where <$Con$> denotes a segment.
@@ 5.1 Settings
Baselines. We incorporate 13 representative LVLMs into the proposed framework including Gemini 1.5 Pro [20], GPT-4V [19], GPT-4o, LLaVA-1.6 [23], XComposer2 [26], Monkey [25], CogVLM [28], MiniCPM-V2 [33], mPLUG-Owl2 [29], Qwen-VL [27], MMAlaya [32], Xtuner [30] and VisualGLM [31]. We implement Gemini 1.5 Pro, GPT-4V and GPT-4o via their official webs and APIs. For other LVLMs, we implement them based on the VLMEvalKit toolkit [34].
Evaluation. We evaluate LVLMs on the test set of our UKD-VQA dataset. In addition to evaluating LVLMs via VLMEvalKit, we design additional matching patterns for each LVLM with respect to its answer format. For example, we additionally use the pattern “The answer is XXX." for XComposer2 as it often answers in this format. All evaluations are conducted with a single Nvidia A100 GPU.
5.2 Quantitative Comparison with SOTA LVLMs
We compare with state-of-the-art LVLMs on the UDK-VQA test set, including Gemini 1.5 Pro [20], GPT-4V [19], GPT-4o and LLaVA-1.6 [23]. For Gemini 1.5 Pro, GPT-4V and GPT-4o, we implement their Raw version via official APIs, which do not have the ability of IAG. Since Gemini 1.5 Pro is famous for its ability to receive long contexts, we use all website content returned by the search engine of our framework to prompt it directly, dubbed LC. For GPT-4V and GPT-4o, we test their self-contained IAG-capable ability via prompting their official web versions with “Retrieve relevant news and answer the question directly from the given options using the option letters based on the image.", dubbed IAG. We incorporate each Raw baseline into our framework as Ours.
Experimental results are listed in Table 1, we can observe that: (1) GPT-4o with our framework achieves the best performance on almost categories of UDK-VQA. (2) For all four baselines, our framework consistently improves their accuracy (e.g., 22.7% and 34.0% absolute performance gains in overall accuracy for GPT-4V and GPT-4o, respectively). (3) Our framework uses shorter contexts but has higher accuracy (e.g., 76.4% vs 83.3% in accuracy for LC and Ours variants of Gemini, respectively). The observations suggest that our framework is generalizable and effective in enhancing the ability of LVLMs to answer questions about up-to-date knowledge.
Table 1: Comparision with SOTA LVLMs on UDK-VQA, where Raw represents the model without IAG ability (e.g., official API version), IAG represents the model with self-contained IAG-capable ability (official web version), LC represents the model with long context input. “Ours" stands for incorporating the Raw baseline into our framework. The value outside/in () indicates the accuracy over samples that do not violate the content management policy of current/all model(s).
@@ 5.2 Quantitative Comparison with SOTA LVLMs
LLaVA Raw 43.8 (45.7) 32.3 (31.8) 22.4 (23.8) 24.9 (23.2) 25.0 (25.0) 53.3 (55.4) 33.3 (31.5) 31.8 (31.2)
1.6 Ours 86.7 (87.6) 91.9 (91.8) 93.9 (97.6) 88.1 (88.0) 90.9 (90.0) 93.3 (93.8) 95.5 (100) 90.2 (90.7)

### v2

# SearchLVLMs: A Plug-and-Play Framework for Augmenting Large Vision-Language Models by Searching Up-to-Date Internet Knowledge

@@ Abstract
Large vision-language models (LVLMs) are ignorant of the up-to-date knowledge, such as LLaVA series, because they cannot be updated frequently due to the large amount of resources required, and therefore fail in many cases. For example, if a LVLM was released on January 2024, and it wouldn’t know the singer of the theme song for the new Detective Conan movie, which wasn’t released until April 2024. To solve the problem, a promising solution motivated by retrieval-augmented generation (RAG) is to provide LVLMs with up-to-date knowledge via internet search during inference, i.e., internet-augmented generation (IAG), which is already integrated in some closed-source commercial LVLMs such as GPT-4V. However, the specific mechanics underpinning them remain a mystery. In this paper, we propose a plug-and-play framework, for augmenting existing LVLMs in handling visual question answering (VQA) about up-to-date knowledge, dubbed SearchLVLMs. A hierarchical filtering model is trained to effectively and efficiently find the most helpful content from the websites returned by a search engine to prompt LVLMs with up-to-date knowledge. To train the model and evaluate our framework’s performance, we propose a pipeline to automatically generate news-related VQA samples to construct a dataset, dubbed UDK-VQA. A multi-model voting mechanism is introduced to label the usefulness of website/content for VQA samples to construct the training set. Experimental results demonstrate the effectiveness of our framework, outperforming GPT-4V by $\sim$25in accuracy.
@@ 1 Introduction
Large vision-language models (LVLMs, e.g., GPT-4V [19], Gemini Series [20], and Grok [21]) have received much attention for their impressive generative capabilities. They require a large resource for data collection, cleaning, and training, restricting them from frequently updating models. However, new information and knowledge are created every time, making LVLMs ineffective in many scenarios. For example, if we talk with LLaVA-1.6 [23] (released on January 30, 2024) about the new Detective Conan movie (realeased on April, 2024), such as “the singer of the theme song”, it performs very badly. It is promising to augment LVLMs by retrieving up-to-date knowledge via internet search during inference, i.e., internet-augmented generation (IAG). Although commercial LVLMs such as GPT-4V [19] and Claude3 [22] have the ability of IAG, the specific mechanics underpinning them remain undisclosed. This paper proposes a plug-and-play framework to augment different LVLMs in handling visual question answering (VQA) about up-to-date knowledge, named SearchLVLMs.
@@
Figure 1: The proposed SearchLVLMs, a framework for LVLMs to access up-to-date knowledge.
@@
To validate the effectiveness and generalizability of the proposed framework, we incorporate 15 state-of-the-art LVLMs into the framework, such as GPT-4V [19] and LLaVA-1.6 [23]. Notably, once the hierarchical filtering model is trained, our framework can adapt different LVLMs and improve their performance without any fine-tuning. Extensive experimental results demonstrate that our framework can significantly improve LVLM’s ability to answer questions about up-to-date knowledge. Incorporating the LLaVA-1.6 model of our framework even outperforms the self-contained IAG-capable GPT-4V by $\sim$25in accuracy on UDK-VQA test set.
Our contributions are summarized as follows. (1) We propose the first open-source framework seamlessly incorporating existing LVLMs with up-to-date knowledge during inference. (2) We propose a pipeline that automatically generates VQA samples related to up-to-date news and construct the first test set for evaluating LVLMs’ ability to handle VQA on up-to-date knowledge. (3) Extensive experimental results on 15 state-of-the-art LVLMs demonstrate the effectiveness of our framework.
@@ 2.2 Large Models with Search Engine
3 SearchLVLMs Framework
In this section, we introduce SearchLVLMs, a framework that seamlessly incorporate existing LVLMs, allowing these LVLMs to access up-to-date knowledge without fine-tuning. The whole framework is illustrated in Figure 1. For a natural language question $Q$ about an image $V$, we first extract queries for both $Q$ and $V$ via the query generator. Then we enter the queries into search engines, and the search engine would return related websites, each of which consist of a title and a snippet. To identify the most helpful content within the websites, a website filter is used to filter the websites based on their titles and snippets, and a content filter is further used to filter the content of the websites filtered by the website filter. Finally, we stitch the filtered content together to prompt existing LVLMs.
@@ 3.1 Query Generator
Question Query Generator. To get queries that make search engines return websites containing helpful content, we leverage large language models (LLMs) to extract queries for $Q$. Thanks to the language understanding capability of LLMs, the role played by each word can be well inferred from the grammatical information of $Q$, even if certain words are unknown for the LLMs. We use “Do not try to answer the question, just print the most informative no more than three entities in the question. Put them on one line and separate them with comm.” to prompt LLMs to generate queries.
@@ 3.2 Search Engine
The extracted queries are fed into a search engine, and the search engine returns relevant websites with their titles and snippets. However, the returned titles and snippets often contain limited and incomplete information. For example, for a website with title “Pororo Dragon Castle Adventure”, the entire snippet returned by Bing is “Pororo and his friends were having fun when a little red dragon named Arthur appears above them Arthur who claims to be the king of dragons commands Pororo and his friends to search for his Dragon …”, obviously there is more about “Pororo Dragon Castle Adventure” contained in the website. Thus we parse the textual content of all websites. For a website, not all of its content contributes to answering questions, we empirically divide the website content into segments every third sentence for a more granular selection of content.
@@ 3.3 Hierarchical Filtering Model
Website Filter. The aim of the website filter is to perform the filtering of websites based on their titles and snippets. Specially, a website scoring model is trained via instruction tuning, to predict how helpful a website will be in answering a question, and the $N$ websites with higher scores would be kept. The training samples are in the format $(T,S,Q,V,R_{w})$, where $R_{w}$ is a quantitative usefulness in the interval $[0,1]$ representing how helpful a website with title $T$ and snippet $S$ will be in answering $Q$ related to $V$. Based on the samples, we construct instructions like “How helpful is an article with such a title and snippet in answering the question based on the image? Choose the best option. Title: <$T$> Snippet: <$S$> Question: <$Q$> Options: A. 1.0 B. 0.8 C. 0.6 D. 0.4 E. 0.2 F. 0.0”. In doing so, the score regression problem is converted into a classification problem, which is easier to learn.
Content Filter. The content filter is used to select the most helpful content segments from the websites filtered by the website filter. For each content segment, we predict how helpful is it for answering $Q$ by a content scoring model. The content scoring model is trained by samples in the format $(C,Q,V,R_{c})$, where $C$ is a content segment, and $R_{c}$ is the quantitative usefulness of $C$ in answering $Q$. The instructions for training the content scoring model are in the format: “How helpful is this context in answering the question based on the image? Choose the best option. Context: <$C$> Question: <$Q$> Options: A. 1.0 B. 0.8 C. 0.6 D. 0.4 E. 0.2 F. 0.0”. We use the model to sort all content segments and select the $M$ highest scoring ones as the obtained segments.
@@ 3.4 Augmented Generation
We augment existing LVLMs by prompting them with the final obtained content, to improve their ability of answering questions about up-to-date knowledge. Taking answer the multiple choice questions as an example, for a question $Q$ with candidate answers $A_{1}$, $A_{2}$, $A_{3}$ and $A_{4}$, we use the prompt “Given context: <$X$> Question: <$Q$> Answers: A.<$A_{1}$> B.<$A_{2}$> C.<$A_{3}$> D.<$A_{4}$> Answer with the option’s letter from the given choices directly based on the context and the image.”, where $X$ denotes the final content obtained by the hierarchical filtering model.
@@ 4.2 Question Generation
For each query, we use Bing to search for relevant and up-to-date news. For the scraped news content, we divide every third sentence into a segment, and use the following message to prompt GPT-3.5 to generate a question-answer pair and several confused answers for each segment: “Given context: <$Con$> Filling the blanks to generate a question about the most informative event of the context, generate an correct answer to the question in no more than three words based on context, and generate three incorrectly confused answers of no more than three words based on context. Question:       Correct answer:       Incorrect answers: A.       B.       C.      ”, where <$Con$> denotes a segment.
@@ 5.1 Settings
Baselines. We incorporate 15 representative LVLMs into the proposed framework including Gemini 1.5 Pro [20], GPT-4V [19], GPT-4o, InternVL-1.5 [43], LLaVA-1.6 [23], LLaVA-1.5 [37] XComposer2 [26], Monkey [25], CogVLM [28], MiniCPM-V2 [33], mPLUG-Owl2 [29], Qwen-VL [27], MMAlaya [32], Xtuner [30] and VisualGLM [31]. We implement Gemini 1.5 Pro, GPT-4V and GPT-4o via their official webs and APIs. We implement other LVLMs based on VLMEvalKit [34].
Evaluation. We evaluate LVLMs on four datasets including GQA [44], InfoSeek [45], A-OKVQA [46] and the proposed UDK-VQA. The reason behind selecting GQA, InfoSeek, and A-OKVQA is to evaluate the generalization capability of our framework across datasets that do not necessitate up-to-date knowledge. In addition to evaluating LVLMs via VLMEvalKit, we design additional matching patterns for each LVLM with respect to its answer format. For example, we additionally use the pattern “The answer is XXX.” for XComposer2 as it often answers in this format. All evaluations are conducted with a single Nvidia A100 GPU.
Table 1: Comparision with SOTA LVLMs on UDK-VQA, where “Raw” represents the model without IAG ability (e.g., official API version), “IAG” represents the model with self-contained IAG-capable ability (official web version), “LC” represents the model with long context input. “Gen.”, “Cham.” and “CLIP→FID (C→F)” denote the method from [49], [13] and [45], respectively. “” indicates that the method leverages our framework to access up-to-date knowledge. “Ours” stands for incorporating the Raw baseline into our framework. The value outside/in () indicates the accuracy over samples that do not violate the content management policy of current/all model(s).
@@
Raw 43.8 (45.7) 32.3 (31.8) 22.4 (23.8) 24.9 (23.2) 25.0 (25.0) 53.3 (55.4) 33.3 (31.5) 31.8 (31.2)
LLaVA Gen. 39.1 (40.0) 31.5 (28.8) 18.4 (19.0) 25.7 (25.3) 36.4 (37.5) 44.0 (44.6) 39.4 (37.0) 31.3 (30.4)
1.6 Cham. 58.6 (58.1) 57.3 (57.5) 53.1 (57.1) 65.3 (67.2) 52.3 (52.5) 72.0 (67.7) 74.2 (72.2) 62.3 (62.7)
C→F 55.5 (56.2) 56.5 (57.9) 34.7 (35.7) 54.0 (54.5) 54.5 (52.5) 62.7 (64.6) 56.1 (53.7) 54.7 (55.3)
Ours 86.7 (87.6) 91.9 (91.8) 93.9 (97.6) 88.1 (88.0) 90.9 (90.0) 93.3 (93.8) 95.5 (100) 90.2 (90.7)
Raw 43.8 (43.8) 53.1 (52.4) 49.0 (47.6) 29.9 (30.7) 34.1 (35.0) 73.3 (76.9) 37.9 (35.2) 42.6 (42.8)
Intern Gen. 29.7 (30.5) 28.1 (26.2) 28.6 (26.2) 22.8 (23.5) 31.8 (32.5) 42.7 (46.2) 28.8 (27.8) 27.6 (27.6)
VL 1.5 Cham. 59.4 (58.1) 61.9 (61.8) 55.1 (57.1) 55.6 (55.4) 52.3 (52.5) 65.3 (67.7) 71.2 (70.4) 59.3 (59.2)
C→F 59.4 (58.1) 65.0 (64.8) 44.9 (42.9) 54.2 (55.7) 47.7 (50.0) 65.3 (66.2) 53.0 (50.0) 57.7 (58.0)
Ours 90.6 (89.5) 95.4 (95.3) 98.0 (97.6) 88.9 (88.0) 100 (100) 96.0 (95.4) 98.5 (98.1) 92.9 (92.3)
5.2 Quantitative Comparison with SOTA LVLMs
We compare with state-of-the-art LVLMs on the UDK-VQA test set, including Gemini 1.5 Pro [20], GPT-4V [19], GPT-4o, LLaVA-1.6 [23] and InternVL-1.5 [43]. For Gemini 1.5 Pro, GPT-4V and GPT-4o, we implement their Raw version via official APIs, which do not have the ability of IAG. Since Gemini 1.5 Pro is famous for receiving long contexts, we use all website content returned by the search engine of our framework to prompt it directly, dubbed LC. For GPT-4V and GPT-4o, we test their self-contained IAG-capable ability via prompting their official web versions with “Retrieve relevant news and answer the question directly from the given options using the option letters based on the image.”, dubbed IAG. We incorporate each Raw baseline into our framework as Ours.
The experimental results on UDK-VQA are listed in Table 1, we can observe that: (1) InternVL-1.5 with our framework achieves the best performance on almost categories of UDK-VQA. (2) For all four baselines, our framework consistently improves their accuracy (e.g., 22.7% and 34.0% absolute performance gains in overall accuracy for GPT-4V and GPT-4o, respectively). (3) Our framework uses shorter contexts but has higher accuracy (e.g., 76.4% vs 83.3% in accuracy for LC and Ours variants of Gemini, respectively). The observations suggest that our framework is generalizable and effective in enhancing the ability of LVLMs to answer questions about up-to-date knowledge.
In addition, the experimental results on GQA, InfoSeek and A-OKVQA are listed in Table 2. Since these datasets do not rely on the up-to-date knowledge, we use a simple strategy to avoid misleading LVLMs with the up-to-date knowledge by invoking our framework when they respond with “E” (as mentioned in Section 4.5) without retrieval. From the table, we can observe that our framework improves the performance of different LVLMs across various datasets. The improvements on these three datasets are not as significant as on our UDK-VQA dataset for the following reasons: (1) The GQA dataset does not rely on external knowledge and is used to evaluate the reasoning ability of LVLMs, which is beyond the scope of our framework. (2) Our framework focuses on retrieving the up-to-date knowledge, whereas the InfoSeek dataset and the A-OKVQA dataset rely on commonsense knowledge, much of which has already been used in the training data of LVLMs.
Table 2: Experiments on GQA [44], InfoSeek [45], A-OKVQA [46], where GQA does Not Rely on external Knowledge (NRK), InfoSeek and A-OKVQA Rely on Commonsense Knowledge (RCK).
Model Variant Retrieval Sources Datasets
Local Data Internet Data GQA InfoSeek A-OKVQA
(clear GT) (unclear GT) (NRK) (RCK) (RCK)
CFR [48] - - - 72.10 - -
Oracle→FID [45] - $\checkmark$ - - 45.60 -
Omni-SMoLA [47] - $\checkmark$ - - - 84.10
LLaVA-1.6 Raw - - 61.66 37.86 75.53
Ours - $\checkmark$ 62.33 41.25 76.22
InternVL-1.5 Raw - - 74.03 51.13 84.53
Ours - $\checkmark$ 74.41 53.10 84.59